Day Shift: 06/15/99

Jason LaDue (jladue@rhodes.colorado.edu)
Tue, 15 Jun 1999 14:56:39 -0600 (MDT)

Shift Summary for Day shift, 06/15/99

On Shift:
Jason LaDue
Miguel Barrio

Other People Present: Ed Blucher, Rick Kessler, Jason Hamm, Marj,
Patty McBride, Peter Shanahan, Elizabeth Turner, Bob T., Hogan,
Leo Bellantoni, Sasha Ledovskoy, Sasha Golossanov, Elliot, and Maxim.

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time:

No Crate 16 FERA ADCs and CsI trouble.

========================================================================
Runs ended this shift:
13603 - Long Laser Scan
13604 - COSMIC, Banana Debugging, No MA ADC crate.
13606 - COSMIC, No MA ADC crate.
13607 - COSMIC, No MA ADC crate.
13608 - COSMIC, mostly No MA ADC crate.
13609 - COSMIC
13610 - KTEV_PED

Tapes brought to FCC this shift: 0

========================================================================
Accesses during shift: None

========================================================================
Description of Shift:
08:40 - Branch 3, Crate 16 was removed from readout to fix power supply.
09:25 - Changed dummy L2 from 50us back to 8us.
10:40 - V' (15) voltage increased by 20V.
10:40 - MA channel 8 is plugged in.
11:50 - Changed the Droges controlling the DC voltages. Done in order
to set the DC1 and DC2 window planes 100 Volts lower to reduce
the number of isolated single hits that have recently become
apparent.
12:00 - Since at least run 16302, Banana 1Y-4 downstream had no
illumination. Fixed by plugging in a loose cable on the
repeater board downstairs.
12:30 - MA DC offsets changed from -10mV to -20mV.
12:30 - Interlocks dropped, DOH!
12:50 - CsI channel 1817 fixed by replacing the Pipeline Module.
13:11 - No spill signal.
13:14 - Run Summary log files have been fixed by a dart fresh start.
13:24 - Spill signal returns.
13:40 - Crate 16 has been swapped.
14:05 - Adjust DC offsets on BA channel 17 - 24 back to dc_offsets to
-5mV. They had drifted at the base.
14:38 - Interlocks have been restored.

run spill time comment
------------------------------------------------------------------------
13603 7:44 Ended Run.
13608 13:11 No Spill Signal
13608 13:24 Spill Signal returns
13608 13:55 Ended Run, Updated CSQ6
13609 14:45 Ended Run, Updated CSQ6 again
13610 15:03 Started Run
13610 15:15 Ended Run, Updated ADGN